A block weighing 48n is pull along a horizontal surface with a rope.The applied force due to the rope is 40n as 37degree to the horizontal.if the block accelerate as 5m/s how large a frictional force must be retarding it motion


1
Expert's answer
2021-09-17T09:22:41-0400

"F_x=40\cdot\cos37°\approx32\ (N)" ;


"m=48/9.81=4.1\ (kg)"


"F=ma\to -F_f+F_x=ma\to F_f=F_x-ma=32-4.1\cdot5=11.5\ (N)" . Answer